Please help me solve this Open-ended problem:
At one point early this season, Clipper was batting .500. After getting 3 hits in 10 at-bats in his next games, his average dropped to .333.
B. How many hits and at-bats must he have altogether.

Answer by edjones(8007)

This is tough especially if you don't follow baseball.
Let x be the number of hits when he batted .500, 2x are the number of at bats.
x/2x=1/2=.500
3/10=.300
.
(x+3)/(2x+10)=1/3
3(x+3)=2x+10 Multiply each side by 3(2x+10) to eliminate fractions.
3x+9=2x+10
x=1
So, he had 1 hit in 2 at bats.
.
Check:
(1+3)/(2+10)=1/3=.333
So he had 4 hits in 12 at bats altogether. Answer.
.
Ed